UT 20140116 LUCI2 Nighttime Commissioning

Observer: JHeidt, APramskiy, MEdwards, DThompson (remote)
Telescope Operator: SAllenson
LBTO Support: TSargent

Plan for the night

- New masks need to come to thermal equilibrium, ie only imaging/LS is possible this night
- Twilight flats
- Get rid of the tilt of the offsets implied by the telescope, requires AGw-transform data collection
- Full set of ZPs in all filters again
- Persistence/crosstalk on sky (perhaps with both LUCIs in parallel?)
- W3 (if seeing is good)
- Image movement due to flexure (if pointing mask is already cold)
- Tests LS-acquisition
- Throughput gratings

- Correct field alignment, check with Alexander

Summary

- Took a set of AGw-transform data

- Took a set of data to check for persistence/crosstalk on sky

- Full set of ZP's with N3.75 camera in all filters

- Determined throughput gratings (some may need to be redone)

- LS-acq. tests

- Generally, GEIRS, SW + RTD are more stable, but
-- memory allocation for GEIRS must be optimized
-- SW-problem with the freezing terminal must be fixed

Preparations

- Alexander/Kai have revised the SW to improve stability of SW and interaction with GEIRS
- Ingo has revised his RTD to make it more reliable and user friendly
- Dave has prepared new MOS-masks on astrometric fields for MOS-acq tests
- A cabinet exchange by Peter with LBTO support is planned for the day preceding the 1st com night

Details

All times are UT, files are luci2.20140115.iiii

00:00 mask exchange during the day was successfull

00:30 Weather is photometric, T=+2deg, Hum=20%, Wind~8-10m/s, team is up on the mountain

00:42 Sunset

01:00 Alexander/Tom were busy in preparing their SW, too late for flats

Point to W_10 to do offset-tests
#5, origin 1160, 1150
#6, pos. detxy Steve 1270,1140
#7, pos detxy Steve, 1455, 1129
#8,9 origin, 1212,1100
#10, pos detxy LUCI telescope service, 1464,1269
#11, origin, 1227, 1067
#11,12 pos detxy offset from RTD, 1341,1045
=> all consistent, flip sign removed

01:20 Steve did pointing & collimation, something is wrong, guider moved GS away from hotspot
instead of putting him into it, DThompson & JHill are working on it

01:30 Alexander uploaded a new SW-version

02:20 Problem fixed, guider works, here is a summary by Dave of what have been done

To orient KFP to the LUCI2 detector:
PCSInstrument.conf RIGHTXYFLIP changed to "true false"

To get the guide probe to move to the correct off-axis location:
oacontrol.conf M_XGP changed to -1

To get guiding and acquisition offsets to go in the correct direction:
LUCI_R.cfg pointingCorr_flip_x changed to +1, GuidingCorr_flip_x changed to +1

To get the wfs applying collimation corrections in the right direction:
LUCI_R.cfg WFS_pupil_dir_x changed to +1

To get LUCI2 talking to the TCS in the same coordinate system:
Remove the sign flip in X from the Telescope Service, recompile and install

Photometric ZP's in all filters

02:31 BS9107, BS9107_FW2, ZPs of all filters in FW2 (JHKKsz),
stopped with "instrument manager" error (instr.manager not initialized),
initialized instr.manager and resent, OK, #13-22
FW2 got stuck when moving to K-filter, recovered, cont'd,
script stopped after 3rd K-exposure without error, #23-25, stopped/restarted scheduler, OK
#26-29 crap, restartet SW, K-band exposures 4/5 and 5/5 #30/31,
probably communication problem with telescope (no response from it)
finished up with Ks and z, #32-41, seeing variable, 0.7-1.2"

03:38 BS9107, BS9107_FW1, ZPs of all filters in FW1 (Br_gam, OH_1060, Y1, Y2
P_gam, P_beta, J_low, J_high, OH_1190, FeII, H2, HeI
Stopped after the 4th Br_gam exp #42-45 (again communication problem with telescope)
stopped/started telescope service, cont'd last Br_gam #46,
cont'd #47-61, stopped again at 1st P_gam, stopped/restart tel.manager & scheduler
#62 crap, cont'd, stopped again at 4th J_high-exp (always offsets are not communicated to telescope),
recovered as usual, cont'd, #47-100

AI Alexander/Tom this bug must be removed

05:27 BS9107_jitterxy, 3 images, center, dy +40", dy -80", check if 2.6deg angle still holds #101-103
quick check by Michelle, confirmed

06:00 Prep for new set of AGw-transform data, chg of RIGHTZEROPOINT

Check tilt again, #105 (center), #106 (dy +40"), #107 (dy-80"), tilt=0.5" over 40"
seems OK, but Michelle is not sure, test it via LS

In the meantime screen is frozen (as in Dec), Alexander/Tom suspect that this is not
due to LUCI-SW but rather due to the NX-client. Will test this later this night

06:40 Steve points to a different (higher) field for collimation, PCS error, slitview #108/109, star at various pos #110-113, subtracted star - LS-ima, seems OK, shift is along slit

AGW-transform data

06:56 Darks for AGw-data, #114/115 (8*4sec integrated)

07:07 Collimate on-axis at StoneE

07:30 Michelle started to take AGw-transform data on StoneE, files #116-145

Persistence/crosstalk

09:43 Preset to FS17, FS17.xml, 30 positions in 20" jitter-box, 6*10sec int, field coveres stars with
a wide range of mags, seeing 0.6-0.8", #146-175 , time required: 48min => Duty Cycle =63%
Alexander started a 2nd NXclient-session in parallel for test-purposes

ZP's in BB-filters 3.75/1.8 camera

10:31 BS9142.xml, ZPs in zJHKKs, N3.75-camera, #176-185

10:43 BS91421_1-8cam.xml, ZPs in zJHKKs, N1.8-camera, #186-195
Aladin bug (complains about change of pixel-scale) still present)

Throughput-gratings

10:52 BS9142_all_corrected.xml, 3"-flexspec, sky #196, src #197, slitview #198, verification #199
applied small offset dx = -0.5", #200, OK
#201-204 210z@0.96, GEIRS error @ first J-exp. (could not allocate memory),
recovered, restartet J-band spec, #205-208 210J@1.25,#209-212 210H@1.65, #213-216 210K@2.2
stopped and realigned star with slit, was off by 11pix in x, no idea why, confirmation #217-227
#228-231 200HK@1.93, better, but without proper AGw-transform (+ hence removal of 2deg tilt) this does not make much sense, should be redone later on,
#232-235 200HK@1.17, #236-239 150Ks@2.17, seeing very variable, 0.6-2"

12:45 RTDv3 uploaded and installed

12:50 Telluric, HIP53735, 3"-flexspec, #240-242 sky, #243-245 src, #246-248 3"-LS, alignment & verification #249-250,OK, #251/252 210z@0.96, #253/254 210J@1.25, didn't want too stabilize at 1.25mue once, redo OK,
#255/256 210H@1.65, #257/258 210K@2.2
#259/260 200HK@1.93, #261/262 200HK@1.17, #263/264 150Ks@2.17, than stop (cannot tilt) <br

Curiously Ks-grating could be tilted properly tonight, 210J@1.25mue failed two times

RTDv3 works even better, several bugs removed

13:20 BS9158, PA=281, LS-acq. with two stars, sky #265, src #266, LS #267
aligned, verification #267-269, rotation OK (stars nicely lined up), but offset in x about 45pix

13:50 Move to zenith for flats

#270-279, J, 10*3sec, 10k
#280-289, J, 10*4sec, 25k
#290-299, H, 10*3sec, 20-30k

Yak, terminal frozen.....Alexander checks it (using the 2nd NXclient)
Nada, this is not the source of a freezing terminal, must be in the LUCI-SW

14:15 Closing...

14:25 Handover to DThompson for calibrations

-- JochenHeidt - 16 Jan 2014
Topic revision: r9 - 16 Jan 2014, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback